OpenStudy (anonymous):

Help plz: Use elimination to solve the system given below. 17x - 2y = 25 17x + 3y = 5 A: (5, 0) B: (1, -4) C: (-1, -5) D: (2, 10)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thank you, Plz give good answer so i can rank once if you could :P

OpenStudy (anonymous):

Nm i will just help some people out

OpenStudy (anonymous):

17x - 2y = 25 17x + 3y = 5 - - - _______________ 0 - 5y = 20 y = -4